# VGS-ATD: Robust Distributed Learning for Multi-Label Medical Image Classification Under Heterogeneous and Imbalanced Conditions

### v1

@@ V-B Phase2: Classification on 15 Nodes
Datasets Classes Samples Central Learning Federated Learning Swarm Learning VGS-ATD Acc Pre Sen F1 Acc Pre Sen F1 Acc Pre Sen F1 Acc Pre Sen F1 Chest X-Ray* 4 21,181 91.89 91.89 91.29 91.59 74.35 74.35 73.85 74.10 68.59 68.59 65.68 67.10 97.06 97.06 97.03 97.04 Brain Cancer 4 5,712 72.62 72.62 76.77 74.64 52.78 52.78 63.49 57.64 39.36 39.36 55.97 46.22 87.49 87.49 87.96 87.72 Brain Tumor 2 3,695 94.86 94.86 87.52 91.04 87.55 87.55 83.59 85.53 82.29 82.27 70.86 76.14 92.96 92.96 92.71 92.84 Breast Cancer 2 2,080** 69.47 69.47 69.81 69.64 62.26 62.26 65.40 63.79 50.00 50.00 58.92 54.10 78.37 78.37 78.55 78.46 Breast Tumor 2 16,826 92.98 92.98 93.01 92.99 63.48 63.49 61.87 62.67 55.55 55.55 54.05 54.79 92.29 92.29 92.21 92.25 Chest Covid 3 13,685 91.01 91.01 91.68 91.35 75.81 75.81 76.65 76.23 64.19 64.19 66.76 65.45 92.44 92.44 82.71 92.57 Fracture Spine 2 3,800 68.25 68.25 68.08 68.16 91.50 91.50 85.51 88.41 76.25 76.25 65.03 70.20 80.25 80.25 80.05 80.15 Fracture Body 2 6,400 89.72 89.72 86.48 88.07 63.83 63.83 63.58 63.71 55.73 55.73 50.90 53.21 98.22 98.22 97.07 97.64 Kidney Cancer 2 8,000 97.56 97.56 96.78 97.17 77.38 77.38 73.73 75.51 70.50 70.50 57.49 63.34 96.44 96.44 95.54 95.99 Liver CT 2 2,846 90.70 90.70 93.49 92.07 57.19 57.19 69.07 62.57 25.96 25.96 64.91 37.09 87.89 87.89 89.95 88.91 Liver Fibrosis 5 5,061 75.59 75.59 75.59 75.59 50.59 50.59 55.17 52.78 34.19 34.19 44.19 38.55 97.73 97.73 97.73 97.73 Lumbar Coordi 3 3,000** 95.02 95.02 88.02 91.39 91.54 91.54 67.90 77.97 85.57 85.57 64.42 73.50 95.02 95.02 90.52 92.72 PCOS Ultra 2 12,315 97.48 97.48 98.36 97.92 92.53 92.53 88.16 90.29 80.39 80.39 73.94 77.03 98.25 98.25 98.61 98.43 Overall:15 Sets 35 112,601 89.96 85.33 84.10 83.41 72.35 68.84 67.57 66.78 72.14 67.81 65.56 64.02 93.96 92.26 92.59 92.15 *: the marked dataset contains three sub-datasets (Chest Pneumonia, Tuberculosis, and Covid) based on Phase 1 combination. **: The marked dataset applied data augmentation techniques to enlarge its training group. The Samples column represents the number of training samples for each dataset. The evaluation metrics used: Acc: Accuracy, Pre: Precision, Sen: Sensitivity, F1: F1 score.
@@
In the second phase, we incorporated 12 additional datasets from various grey-scale imaging modalities to simulate a scenario where new datasets are continually introduced into the system. This expansion increases the severity of data heterogeneity and class imbalance. The resulting dataset collection comprises 15 distinct subsets, including three sets inherited from Phase 1, which cover a total of 35 independent class labels. To address these new challenges, we applied the VGS-ATD using its vertical configuration. To conserve computational resources and improve training efficiency, we reused the pretrained weights derived from the Phase 1 CXR dataset as the foundational source domain. The CXR pretrained weight is based on relevant grey-scale features, providing a useful knowledge base for the newly introduced data.
@@
Table III presents a comprehensive performance comparison, showing both the overall classification performance across all 35 classes and the individual performance of the final global model on each dataset. As expected, all learning methods experienced a performance drop compared to phase 1 due to the increased complexity of the expanded dataset. However, the performance degradation observed in the centralized learning and VGS-ATD is significantly less severe than that of federated and swarm learning, highlighting their superior ability to manage heterogeneous data.
@@
Notably, VGS-ATD achieved the best performance across nine datasets, including Chest X-ray, Brain Cancer, and Brain Tumor, while centralized learning led in five datasets. In contrast, federated learning achieved top performance in only one dataset, and swarm learning consistently lagged behind the other approaches. These results suggest that VGS-ATD delivers more stable and balanced performance across datasets, which is highly valuable in real-world medical applications. For instance, although centralized learning reported an overall accuracy of 90%, its performance deteriorated on small datasets such as Breast Cancer and Fracture Spine, due to its susceptibility to class imbalance.
@@
Another key advantage of VGS-ATD is its resilience to catastrophic forgetting. When comparing performance on the Chest X-ray dataset between the two phases, all baseline methods showed a drop in accuracy. This is due to the phenomenon known as catastrophic forgetting, where central and federated learning methods require retraining the entire global model on new data, resulting in the model forgetting its previous knowledge. This issue arises because these baselines integrate all knowledge into a central model from scratch each time a new node is introduced. The reason for performance decreasing in swarm learning is different from the other two methods, mainly due to its sequential learning paradigm uses the aggregated knowledge after each round of training to initialise the next round’s base node, and causes catastrophic forgetting when each round finishes. In contrast, our method preserves the previously trained weights, enabling it to retain and reuse knowledge learned in earlier phases. This ensures performance consistency across evolving data environments.
Figure 6: Phase2 training time and computational resource cost comparison.
@@
Fig. 6 further compares the training time and computational resource consumption across the four methods. Federated learning and VGS-ATD demonstrate a clear advantage in terms of time efficiency over centralized and swarm learning. As the dataset size grows, methods using a parallel learning setting exhibit a greater advantage over those that use a sequential learning setting.
@@
V-C Phase3: Classification on 29 Nodes
TABLE IV: Phase3 performance comparison on 29-node classification task.
@@ V-C Phase3: Classification on 29 Nodes
Datasets Classes Samples Central Learning Swarm Learning VGS-ATD Acc Pre Sen F1 Acc Pre Sen F1 Acc Pre Sen F1 Chest X-Ray* 4 21,181 75.98 75.98 75.56 75.77 73.61 73.59 72.45 73.02 96.17 96.20 96.20 96.20 Brain Cancer 4 5,712 78.41 78.43 79.44 78.93 54.00 54.01 67.11 59.85 85.96 85.96 86.76 86.36 Brain Tumor 2 3,695 93.91 93.91 92.66 93.28 78.89 78.90 75.32 77.07 92.96 92.96 93.60 93.28 Breast Cancer 2 2,080** 50.48 50.48 50.72 50.60 61.29 69.08 62.20 61.74 76.44 76.44 77.18 76.81 Breast Tumor 2 16,826 94.06 94.06 94.04 94.05 60.07 61.29 58.01 59.02 89.62 89.63 89.55 89.59 Chest Covid 3 13,685 85.12 89.80 91.92 90.85 69.09 69.08 72.48 70.74 91.56 90.77 91.63 91.20 Fracture Spine 2 3,800 66.00 65.99 62.56 64.23 68.50 68.49 63.57 65.94 69.75 69.75 69.75 69.75 Fracture Body 2 6,400 78.85 78.85 81.93 80.36 59.09 59.08 56.63 57.83 98.61 96.79 96.71 96.75 Kidney Cancer 2 8,000 96.87 96.87 97.98 97.42 74.50 74.50 65.32 69.61 95.69 95.70 95.03 95.36 Liver CT 2 2,846 89.64 89.66 86.46 88.03 41.40 41.41 66.11 50.92 90.53 90.53 90.53 90.53 Liver Fibrosis 5 5,061 79.25 79.25 79.17 79.21 47.04 47.04 53.24 49.95 97.92 97.92 97.54 97.73 Lumbar Coordi 3 3,000** 97.51 97.52 93.33 95.38 91.04 94.94 69.32 91.04 96.52 96.52 90.23 93.27 PCOS Ultra 2 12,315 91.82 94.39 92.78 93.58 91.64 91.63 86.44 88.96 97.60 97.61 98.24 97.92 ALL Cancer 4 16,000 97.25 97.15 96.85 97.00 93.28 93.28 91.20 93.28 98.66 98.66 98.38 98.52 Breast Cell 2 8,000 84.43 84.44 90.13 87.19 81.19 81.19 80.73 80.96 92.63 92.63 93.62 93.12 Blood Cell 5 16,000 89.88 89.88 89.76 89.82 86.88 87.24 84.81 86.01 93.88 93.88 93.76 93.82 Cervix Cancer 5 20,000 88.87 88.88 88.72 88.80 76.30 76.31 77.44 76.87 95.75 95.92 96.02 95.97 Colon Cancer 5 20,000 90.55 92.37 91.36 91.86 88.33 84.38 79.43 81.83 94.75 95.01 94.51 94.76 Colon Disease 4 12,800 69.75 69.75 70.45 70.10 69.13 69.14 64.30 66.63 78.12 78.12 77.26 77.69 Diabetic Foot 2 2,000** 85.79 85.80 62.50 72.32 40.24 40.24 70.10 51.13 81.07 81.07 90.13 85.36 Lymph Cancer 3 12,000 72.54 72.54 72.18 72.36 61.04 61.03 59.82 60.42 86.54 86.55 86.43 86.49 Melanoma Cancer 2 11,879 71.20 71.21 74.05 72.60 76.35 76.34 73.59 74.94 91.40 91.41 91.17 91.29 Oral Cancer 2 8,002 75.25 75.24 74.88 75.06 54.94 54.93 60.08 57.39 80.81 86.55 86.43 86.49 Otoscopic 5 2,400 74.16 74.17 74.01 74.09 59.79 59.80 69.32 64.21 96.46 96.47 95.46 95.96 Prostate Cell 2 1,000** 81.82 79.42 54.00 64.29 24.24 23.53 47.06 31.37 88.79 96.47 95.46 95.96 Retinal Cataract 2 2,000** 57.69 34.76 71.43 46.76 63.80 68.16 55.58 61.23 87.60 87.60 87.60 87.60 Retinal Glaucoma 2 4,800 65.94 85.39 57.93 69.03 46.56 44.05 65.99 52.83 97.19 97.19 97.09 97.14 Overall: 29 Sets 80 244,128 84.92 82.48 81.27 78.96 72.99 70.91 66.9 66.19 92.64 90.89 91.35 90.61 *: the marked dataset contains three sub-datasets (Chest Pneumonia, Tuberculosis, and Covid) based on Phase 1 combination. **: the marked dataset applied data augmentation techniques to enlarge its training group. The evaluation metrics used: Acc: Accuracy, Pre: Precision, Sen: Sensitivity, F1: F1 score.
@@
In the third phase, we further expanded the scope of target data by introducing a new group of color medical datasets. This second dataset group consists of 14 datasets encompassing 45 independent class labels, including pathology, fundus, endoscopy, otoscopy, and general dermatological images. The total datasets and corresponding labels in this phase reached 29 sets and 80 classes. This phase simulates a real-world scenario in which multiple independently trained dataset groups, previously using VGS-ATD, aim to be integrated to build a large platform. This scenario enables us to evaluate the performance of the ATD method in an ATD-to-ATD configuration.
@@
In this phase, we directly applied the global model obtained from Phase 2 and performed an ensemble with a new global model trained on the incoming dataset group. Notably, compared to the earlier phases, both the sample size and the number of client nodes increased substantially, placing a heavy computational burden on our system. The centralized learning method struggled with longer data loading and computation times. In contrast, distributed learning methods, such as swarm learning and VGS-ATD, remained unaffected, thanks to their ability to train each local model independently before aggregation. However, federated learning became impractical in this setting, as our hardware could not support training and aggregating 29 local models simultaneously. Consequently, only centralized learning and swarm learning were included as baselines in this phase for comparative evaluation.
@@
Another key distinction in this phase involves the choice of source domain for pretraining. Continuing to use the grey-scale CXR dataset as the source domain would likely result in a domain mismatch when applied to the newly introduced colour datasets. To address this, we utilized ImageNet pretrained weights for the new colour datasets, as they provide a broader, general knowledge that includes features relevant to the human body, such as skin and eye structures. This solution enabled faster convergence and more relevant feature extraction. However, since centralized learning requires training a single unified model, it cannot accommodate two different pretrained sources. Thus, we continued to use the CXR dataset as the source domain for the centralized learning baseline, as it performed well in Phase 2. The ImageNet weights were used exclusively with swarm learning and VGS-ATD.
@@
Table IV reveals a consistent trend with Phase 2, showing VGS-ATD outperforming all baselines in terms of overall classification performance. As the dataset size and complexity increased, the advantages of VGS-ATD became even more pronounced. It delivered the best performance across 22 datasets and achieved comparable results to centralized learning on the remaining ones. Meanwhile, swarm learning continued to underperform significantly. Importantly, the Chest X-ray dataset maintained stable and consistent performance with VGS-ATD across all three phases. In contrast, the accuracy of other baselines deteriorated sharply, highlighting VGS-ATD’s resilience to catastrophic forgetting.
Figure 7: Phase3 training time and computational resource cost comparison.
@@
V-D Robustness Analysis
To evaluate the robustness and reliability of VGS-ATD, we performed model attention analysis using Grad-CAM [34], which is a widely used Explainable AI technique. Grad-CAM integrates global average pooling to highlight the critical regions that contribute to model predictions, providing insight into decision-making. In this study, we applied Grad-CAM to the models trained by each method across all three phases of the experiment. Fig. 8 presents representative Grad-CAM visualization samples.
Figure 8: Grad-CAM samples of VGS-ATD and baselines. The title representation includes GT: Ground Truth, CL: Centralized Learning, FL: Federated Learning, SL: Swarm Learning, VGS-ATD: Proposed method. The label representation includes P1-3: Phase 1-3, T: True for the predicted label, and N: Negative for the predicted label.

### v3

@@ 5.2 Phase2: Classification on 15 Nodes
Datasets Classes Samples Central Learning FedLD NonIID-SL VGS-ATD Acc Pre Sen F1 Acc Pre Sen F1 Acc Pre Sen F1 Acc Pre Sen F1 Chest X-Ray* 4 21,181 91.89 91.89 91.29 91.59 75.84 75.84 74.62 75.22 77.00 77.00 75.74 76.36 97.06 97.06 97.03 97.04 Brain Cancer 4 5,712 72.62 72.62 76.77 74.64 58.12 58.12 68.59 62.92 53.17 53.17 65.38 58.65 87.49 87.49 87.96 87.72 Brain Tumor 2 3,695 94.86 94.86 87.52 91.04 85.66 85.66 80.95 83.23 86.06 86.04 80.51 83.19 92.96 92.96 92.71 92.84 Breast Cancer 2 2,080** 69.47 69.47 69.81 69.64 59.62 59.64 61.39 60.49 58.34 57.21 60.10 58.62 78.37 78.37 78.55 78.46 Breast Tumor 2 16,826 92.98 92.98 93.01 92.99 65.46 65.46 63.18 64.30 62.74 62.74 60.49 61.59 92.29 92.29 92.21 92.25 Chest Covid 3 13,685 91.01 91.01 91.68 91.35 74.28 74.28 75.83 75.05 72.05 72.05 72.29 72.17 92.44 92.44 82.71 92.57 Fracture Spine 2 3,800 68.25 68.25 68.08 68.16 74.25 74.25 70.21 72.17 86.26 86.26 81.95 84.04 80.25 80.25 80.05 80.15 Fracture Body 2 6,400 89.72 89.72 86.48 88.07 64.82 64.82 68.33 66.53 60.67 60.67 62.02 61.34 98.22 98.22 97.07 97.64 Kidney Cancer 2 8,000 97.56 97.56 96.78 97.17 75.81 78.94 70.89 73.27 78.94 78.94 70.21 74.32 96.44 96.44 95.54 95.99 Liver CT 2 2,846 90.70 90.70 93.49 92.07 54.56 54.56 69.27 61.04 43.16 43.16 63.90 51.52 87.89 87.89 89.95 88.91 Liver Fibrosis 5 5,061 75.59 75.59 75.59 75.59 53.85 53.85 61.51 57.43 46.94 46.94 53.43 49.97 97.73 97.73 97.73 97.73 Lumbar Coordi 3 3,000** 95.02 95.02 88.02 91.39 95.52 95.52 75.59 84.40 89.05 89.05 71.89 79.56 95.02 95.02 90.52 92.72 PCOS Ultra 2 12,315 97.48 97.48 98.36 97.92 93.06 93.06 88.53 90.74 91.23 91.23 88.22 89.70 98.25 98.25 98.61 98.43 Overall:15 Sets 35 112,601 89.96 85.33 84.10 83.41 72.74 70.22 67.98 67.89 72.14 67.81 65.56 64.02 93.96 92.26 92.59 92.15 *: the marked dataset contains three sub-datasets (Chest Pneumonia, Tuberculosis, and Covid) based on Phase 1 combination. **: The marked dataset applied data augmentation techniques to enlarge its training group. The Samples column represents the number of training samples for each dataset. The evaluation metrics used: Acc: Accuracy, Pre: Precision, Sen: Sensitivity, F1: F1 score.
@@
In the second phase, we incorporated 12 additional datasets from various grey-scale imaging modalities to simulate a scenario where new datasets are continually introduced into the system. This expansion increases the severity of data heterogeneity and class imbalance. The resulting dataset collection comprises 15 distinct subsets, including three sets inherited from Phase 1, which cover a total of 35 independent class labels. To address these new challenges, we applied the VGS-ATD using its vertical configuration. To conserve computational resources and improve training efficiency, we reused the pretrained weights derived from the Phase 1 source dataset, since the pretrained weight is based on relevant grey-scale features, providing a useful knowledge base for the newly introduced data.
@@
Table 3 presents a comprehensive performance comparison, showing both the overall classification performance across all 35 classes and the individual performance of the final global model on each dataset. As expected, all learning methods experienced a performance drop compared to phase 1 due to the increased complexity of the expanded dataset. However, the performance degradation observed in the centralized learning and VGS-ATD is significantly less severe than that of federated and swarm learning settings, highlighting their superior ability to manage heterogeneous data.
@@
Notably, VGS-ATD achieved the best accuracy performance across seven datasets, and demonstrated comparable performance to the leading methods on the remaining datasets, while centralized learning achieved a leading position in five datasets. In contrast, both federated and swarm learning baselines achieved top performance in only one dataset. These results suggest that VGS-ATD delivers more stable and balanced performance across datasets, which is highly valuable in real-world medical applications. For instance, although centralized learning reported an overall accuracy of 90%, its performance deteriorated on small datasets such as Breast Cancer and Fracture Spine, due to its susceptibility to class imbalance.
Figure 6: P2 training time and computational cost comparison.
@@
Another key advantage of VGS-ATD is its resilience to catastrophic forgetting. When comparing performance on the initial three Chest X-ray datasets between the two phases, all baseline methods showed a drop in accuracy. This is due to the phenomenon known as catastrophic forgetting, where central and federated learning methods require retraining the entire global model on new data, resulting in the model forgetting its previous knowledge. This issue arises because these baselines integrate all knowledge into a central model from scratch each time a new node is introduced. The reason for the performance decrease in swarm learning is different from the other two methods, mainly due to its sequential learning paradigm uses the aggregated knowledge after each round of training to initialize the next round’s base node, and causes catastrophic forgetting when each round finishes. In contrast, our method preserves the previously trained weights, enabling it to retain and reuse knowledge learned in earlier phases. This ensures performance consistency across evolving data environments.
@@
Fig. 6 further compares the training time and computational resource consumption across the four methods. FedLD and VGS-ATD demonstrate a clear advantage in terms of time efficiency over centralized and swarm learning. As the dataset size grows, methods using a parallel learning setting exhibit a greater advantage over those that use a sequential learning setting.
@@
5.3 Phase3: Classification on 29 Nodes
Table 4: Phase3 performance comparison on 29-node classification task.
@@ 5.3 Phase3: Classification on 29 Nodes
Datasets Classes Samples Central Learning Non-IID SL VGS-ATD Acc Pre Sen F1 Acc Pre Sen F1 Acc Pre Sen F1 Chest X-Ray* 4 21,181 75.98 75.98 75.56 75.77 75.04 75.04 74.09 74.56 96.17 96.20 96.20 96.20 Brain Cancer 4 5,712 78.41 78.43 79.44 78.93 51.41 51.41 62.87 56.57 85.96 85.96 86.76 86.36 Brain Tumor 2 3,695 93.91 93.91 92.66 93.28 84.84 84.84 82.39 83.60 92.96 92.96 93.60 93.28 Breast Cancer 2 2,080** 50.48 50.48 50.72 50.60 54.09 54.09 55.42 54.74 76.44 76.44 77.18 76.81 Breast Tumor 2 16,826 94.06 94.06 94.04 94.05 60.60 60.60 58.04 59.29 89.62 89.63 89.55 89.59 Chest Covid 3 13,685 85.12 89.80 91.92 90.85 66.35 66.35 68.81 67.56 91.56 90.77 91.63 91.20 Fracture Spine 2 3,800 66.00 65.99 62.56 68.49 82.75 82.75 77.70 80.15 69.75 69.75 69.75 69.75 Fracture Body 2 6,400 78.85 78.85 81.93 80.36 59.09 58.10 65.19 61.44 98.61 96.79 96.71 96.75 Kidney Cancer 2 8,000 96.87 96.87 97.98 97.42 73.06 73.06 67.77 70.32 95.69 95.70 95.03 95.36 Liver CT 2 2,846 89.64 89.66 86.46 88.03 48.94 48.95 64.58 55.69 90.53 90.53 90.53 90.53 Liver Fibrosis 5 5,061 79.25 79.25 79.17 79.21 47.04 47.23 54.69 50.69 97.92 97.92 97.54 97.73 Lumbar Coordi 3 3,000** 97.51 97.52 93.33 95.38 94.03 94.03 63.64 75.90 96.52 96.52 90.23 93.27 PCOS Ultra 2 12,315 91.82 94.39 92.78 93.58 91.64 92.25 85.45 88.72 97.60 97.61 98.24 97.92 ALL Cancer 4 16,000 97.25 97.15 96.85 97.00 92.06 92.06 89.52 90.77 98.66 98.66 98.38 98.52 Breast Cell 2 8,000 84.43 84.44 90.13 82.80 82.81 82.81 80.65 81.71 92.63 92.63 93.62 93.12 Blood Cell 5 16,000 89.88 89.88 89.76 89.82 90.50 90.50 87.86 89.16 93.88 93.88 93.76 93.82 Cervix Cancer 5 20,000 88.87 88.88 88.72 88.80 77.48 77.48 78.69 78.08 95.75 95.92 96.02 95.97 Colon Cancer 5 20,000 90.55 92.37 91.36 91.86 87.01 87.00 84.43 85.69 94.75 95.01 94.51 94.76 Colon Disease 4 12,800 69.75 69.75 70.45 70.10 71.50 71.50 63.49 67.25 78.12 78.12 77.26 77.69 Diabetic Foot 2 2,000** 85.79 85.80 62.50 72.32 49.70 49.70 67.20 57.14 81.07 81.07 90.13 85.36 Lymph Cancer 3 12,000 72.54 72.54 72.18 72.36 56.54 56.54 55.55 56.04 86.54 86.55 86.43 86.49 Melanoma Cancer 2 11,879 71.20 71.21 74.05 72.60 78.25 78.25 78.45 78.35 91.40 91.41 91.17 91.29 Oral Cancer 2 8,002 75.25 75.24 74.88 75.06 60.69 60.69 67.43 63.88 80.81 86.55 86.43 86.49 Otoscopic 5 2,400 74.16 74.17 74.01 74.09 72.50 72.50 76.32 74.36 96.46 96.47 95.46 95.96 Prostate Cell 2 1,000** 81.82 79.42 54.00 64.29 15.15 14.71 83.33 25.00 88.79 96.47 95.46 95.96 Retinal Cataract 2 2,000** 57.69 34.76 71.43 46.76 52.07 56.28 61.98 58.99 87.60 87.60 87.60 87.60 Retinal Glaucoma 2 4,800 65.94 85.39 57.93 69.03 45.63 56.18 63.13 59.45 97.19 97.19 97.09 97.14 Overall: 29 Sets 80 244,128 84.92 82.48 81.27 78.96 73.45 71.46 68.39 67.58 92.64 90.89 91.35 90.61 *: the marked dataset contains three sub-datasets (Chest Pneumonia, Tuberculosis, and Covid) based on Phase 1 combination. **: the marked dataset applied data augmentation techniques to enlarge its training group. The evaluation metrics used: Acc: Accuracy, Pre: Precision, Sen: Sensitivity, F1: F1 score.
@@
In the third phase, we further expanded the scope of target data by introducing a new group of color medical datasets. This color dataset group consists of 14 datasets encompassing 45 independent class labels, including pathology, fundus, endoscopy, otoscopy, and general dermatological images. The total datasets and corresponding labels in this phase reached 29 sets and 80 classes. This phase simulates a real-world scenario in which multiple independently trained dataset groups, previously using VGS-ATD, aim to be integrated to build a large platform. This scenario enables us to evaluate the performance of the ATD method in an ATD-to-ATD configuration.
@@
In this phase, we directly applied the global model obtained from Phase 2 and performed an ensemble with a new global model trained on the incoming dataset group. Notably, compared to the earlier phases, both the sample size and the number of client nodes increased substantially, placing a heavy computational burden on our system. The centralized learning method struggled with longer data loading and computation times. In contrast, distributed learning methods, such as NonIID SL and VGS-ATD, remained unaffected, thanks to their ability to train each local model independently before aggregation. However, FedLD became impractical in this setting, as it takes significant GPU memory to run 29 independent models at the same time, and our hardware setting does not support training them simultaneously. Consequently, only centralized and swarm learning settings were included as baselines in this phase for comparative evaluation.
@@
Another key distinction in this phase involves the choice of source domain for pretraining. Continuing to use the initial source dataset with grey-scale images would likely result in a domain mismatch when applied to the newly introduced color datasets. To address this, we utilized ImageNet pretrained weights for the new color datasets, as they provide a broader, general knowledge that includes features relevant to the human body, such as skin and eye structures. This solution enabled faster convergence and more relevant feature extraction. However, since centralized learning requires training a single unified model, it cannot accommodate two different pretrained sources. Thus, we continued to use the same source domain for the centralized learning baseline, as it performed well in Phase 2. The ImageNet weights were used exclusively with NonIID SL and VGS-ATD.
@@
Table 4 reveals a consistent trend with Phase 2, showing VGS-ATD outperforming all baselines in terms of overall classification performance. As the dataset size and complexity increased, the advantages of VGS-ATD became even more pronounced. It delivered the best performance across 21 datasets and achieved comparable results to centralized learning on the remaining ones. Meanwhile, the performance of NonIID SL does not degrade after node expansion, showing the effectiveness of its design against data heterogeneity. Importantly, the initial three Chest X-ray datasets maintained stable and consistent performance with VGS-ATD across all three phases. In contrast, its performance deteriorated sharply in the other two baselines, highlighting VGS-ATD’s resilience to catastrophic forgetting.
Figure 7: P3 training time and computational cost comparison.
@@
5.4 Robustness Analysis
To evaluate the robustness and reliability of VGS-ATD, we performed model attention analysis using Grad-CAM [zhou2016learning], which is a widely used Explainable AI technique. Grad-CAM integrates global average pooling to highlight the critical regions that contribute to model predictions, providing insight into decision-making. In this study, we applied Grad-CAM to the models trained by each method across all three phases of the experiment. Fig. 8 presents representative Grad-CAM visualization samples.
Figure 8: Grad-CAM samples of VGS-ATD and baselines. The title representation includes GT: Ground Truth, CL: Centralized Learning, and the other three baselines introduced above. The label representation includes P1-3: Phase 1-3, T: True for the predicted label, and N: Negative for the predicted label.